OpenStudy (anonymous):

Melanie has a piece of colth 3 5/6 yd long. How many 13/ yd pieces can be cut from the cloth?

OpenStudy (anonymous):

what size piece is she trying to cut? looks incomplete in your question

OpenStudy (anonymous):

i ment 1/3 yd pieces

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{23}{6}\div \frac{1}{3}=\frac{23}{6}\times \frac{3}{1}=11.5\]

OpenStudy (anonymous):

I converted your mixed number to it's fraction for in the first step

OpenStudy (anonymous):

If the answer is a mixed fraction it should be 11 1/2